OpenStudy (anonymous):

How to factor (x-1)+(x-1)^2

OpenStudy (anonymous):

(x-1)+(x-1)^2= x-1+x^2-2x+1= x^2-x= x(x-1)

OpenStudy (anonymous):

Oh, duh. Thanks!

OpenStudy (anonymous):

just do (x-1)(1+x-1) in the second step and yield (x-1)(x) in the third and final step

OpenStudy (anonymous):

factor did not say solve; correct? (x-1)(1+(x-1)) I agree with the final solved answer but I do not believe that was the question to simplify the answer.

OpenStudy (anonymous):

maybe it was the part of a problem!